Using the information in the table, total fixed costs equals $

Supporting Materials

  • Corn production costs ATC MC Y com TVC TFC TC AVC AFC $10.50 $5.20 $25 10 25 50 70 $2.57 95 100 101 $2.55 $2.77

Question 15

According to the information in the table, what is total variable cost when producing zero bushels of corn?
$

Supporting Materials

  • Corn production costs ATC MC Y com TVC TFC TC AVC AFC $10.50 $5.20 $25 10 25 50 70 $2.57 95 100 101 $2.55 $2.77

Question 16

Costs 4 mgd plant 8 mgd plant 12 mgd plant
Water Intake $19,000,000 $21,000,000 $22,000,000
Water Treatment Plant $41,000,000 $55,000,000 $68,000,000
Design $15,000,000 $18,000,000 $21,000,000
TOTAL COSTS $75,000,000 $94,000,000 $111,000,000

The table below shows the costs to construct a municipal water treatment plant. The table shows 3 possible plant sizes, 4mgd, 8mgd, and 12 mgd (mgd=million gallons a day). Once built, the plant will be used by treat potable water for use by households and businesses.   Based on the data, do water treatment plants exhibit economies of scale?

Yes
No
Cannot calculate from the available data
Answer #1

14) Total fixed cost is $80.

Here AFC = ATC - AVC

At 10 units , AFC = 10.50 - 2.50 [2.50=(25/10)]

AFC = 8

TFC = AFC x Q = 8 x 10 = $80.

15) Total variable cost when producing zero bushels is $0 as only fixed cost is there at 0 units of output.

16) Yes, the water treatment plant does show economies of scale as the additional cost of production is decreasing as the capacity is increased. The rate of increase in cost is falling.